# CT-UIO: Continuous-Time UWB-Inertial-Odometer Localization Using Non-Uniform B-spline with Fewer Anchors

### v1

@@ IV-B Exhibition Hall Dataset
Fig. 7 shows the distribution of control points in sequence B-H2. Fig. 8 shows the overview of trajectories in the exhibition hall dataset. The APE results of the compared methods on the six sequences are summarized in Table IV.
TABLE IV: Performance comparison in APE (RMSE, Meter) of different methods on the exhibition hall dataset. The best results are marked in bold.
Seq Optimization +Filtering[26] Optimization+ Trust Region[27] SFUISE[17] Ours
B-S1 0.432 0.393 0.042 0.040
B-S2 0.372 0.405 0.127 0.122
B-J1 0.760 0.588 0.243 0.219
B-J2 0.391 0.562 0.168 0.098
B-H1 0.885 0.543 0.203 0.150
B-H2 0.728 0.413 0.358 0.311
(a) B-S1
(b) B-S2
(c) B-J1
(d) B-J2
(e) B-h1
(f) B-h2
Figure 8: We align the obtained continuous-time trajectory for all sequences in the exhibition hall dataset with the ground-truth trajectories at the same frequency. The absolute translation error is mapped onto the estimated trajectory for each sequence. It can be observed that our CT-UIO achieves high accuracy in trajectory estimation, closely matching the ground-truth trajectories across all sequences, even in those involving fast turns.
@@
The results notably indicate that CT-UIO achieves an overall improvement in localization performance, even in sequences involving fast turning, thereby verifying the effectiveness of its dynamic knot span adjustment strategy.
@@
Furthermore, it is noteworthy that the discrete-time methods (Optimization+Filtering [26] and Optimization+Trust Region [27]) lead to an increase of APE or even exhibit obvious location deviations when the robot undergoes rapid rotational motion. Compared with the uniform B-splines method (SFUISE), CT-UIO ensures a better capture of motion characteristics by adjusting the control point density. Specifically, CT-UIO reduces the density of control points in sections with less speed variation and increases it during significant speed changes. SFUISE cannot adapt to trajectory complexity, as it sets the knot span as a constant value throughout the trajectory. Besides, in the case of only two UWB anchors, UWB positioning suffers a notable accuracy degradation. To address the observability issue in a few-anchor system, CT-UIO introduces virtual anchors by coupling UWB-ranging data and motion priors from the IMU/odometer fusion model to provide more comprehensive constraints. Based on this, the proposed CT-UIO exhibits robustness performance through multiple virtual anchor ranging constraints even with limited anchor availability.
@@
The results of these experiments in corridor and exhibition hall datasets prove that the proposed CT-UIO is highly effective in situations involving fast translational and rotational motion. Across all sequences, the CT-UIO system maintains more stable and consistent accuracy by utilizing a non-uniform continuous-time trajectory representation. This strategy effectively adjusts the control point interval and ensures better capture of motion pattern characteristics.
V Conclusion
@@
In this article, we present the CT-UIO, a continuous-time UWB-Inertial-odometer localization system designed to achieve accurate and robust localization under few-anchor conditions. CT-UIO tightly integrates UWB ranging, IMU, and odometer data using non-uniform B-spline trajectory representation. Specifically, the proposed system enhances state estimation accuracy in fast-moving scenarios by adaptively placing control points according to varying motion speeds. To ensure the UWB-based localization system is fully observable, CT-UIO combines UWB ranging with motion prior from the IMU/odometer fusion model to generate virtual anchors. Additionally, CT-UIO’s backend conducts adaptive sliding window factor pose graph optimization, incorporating constraints from virtual UWB anchor ranging, IMU, and odometer measurements. The proposed CT-UIO system is evaluated in several real-world few-anchor situations using self-collected datasets. Compared to state-of-the-art discrete-time and continuous-time estimation methods, CT-UIO consistently produces more consistent and robust trajectory estimation, particularly in fast-motion scenes.
@@
In the future, it is worthwhile exploring the implementation of the proposed CT-UIO localization system with unknown anchor positions, as well as incorporating complementary sensors, such as cameras, to enhance the accuracy of short-term pose estimates.

### v2

@@ B-S2 B-J1 B-J2 B-H1 B-H2
Fig. 8 shows the overview of trajectories in the exhibition hall dataset. The APE results of the compared methods on the exhibition hall dataset are summarized in Table VI. The results notably indicate that CT-UIO achieves an overall improvement in localization performance, even in sequences involving fast turning, thereby verifying the effectiveness of its dynamic knot span adjustment strategy.
We also obtain the CDF plots for each sequence, as shown in Fig. 9. It can be seen that the CDF curve of our method is located at the top of the remaining five methods. For sequence B-H2, the proposed method achieves a $90\%$ error of $0.474m$, outperforming Optimization+Filtering ($1.092m$), Optimization+Trust Region ($0.604m$), GPDA ($0.562m$), VMHE ($0.558m$), and SFUISE ($0.536m$). Fig. 10 shows the distribution of control points in sequence B-H2.
@@
Furthermore, it is noteworthy that the discrete-time methods (Optimization+Filtering [yang2023novel], Optimization+Trust Region [zhou2024optimization], GPDA[WOS:001173317800046], and VMHE[9829196] ) lead to an increase of APE or even exhibit obvious location deviations when the robot undergoes rapid rotational motion. Compared with the uniform B-splines method (SFUISE), CT-UIO ensures a better capture of motion characteristics by adjusting the control point density. Specifically, CT-UIO reduces the density of control points in sections with less speed variation and increases it during significant speed changes. SFUISE cannot adapt to trajectory complexity, as it sets the knot span as a constant value throughout the trajectory. Besides, in the case of only two UWB anchors, UWB positioning suffers a notable accuracy degradation. To address the observability issue in a few-anchor system, CT-UIO introduces VAs by coupling UWB-ranging data and motion priors from the IMU/odometer fusion model to provide more comprehensive constraints. Based on this, the proposed CT-UIO exhibits robust performance through multiple VAs’ ranging constraints even with limited anchor availability.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 12: The distribution of control points in sequence C-H1.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 13: CDFs of APE for different methods on the office dataset. The proposed method exhibits the closest performance to the SFUISE in slow sequences while outperforming all compared methods in fast and hybrid sequences, indicating superior accuracy and stability.
V-D Evaluation on the Office Dataset
Lastly, the proposed CT-UIO is evaluated in the office environment of meeting rooms and corridors, using fewer UWB anchors. The robot starts in the corridor, passes through the meeting room, and then returns to the corridor. While two UWB anchors are available at positions $(10.20,3.94)m$, (6.23,2.33)m(6.23,{-}2.33)m in the meeting room $(6.9m{\times}10.5m)$, only a single UWB anchor is available at the position (0.51,0.18)m(-0.51,0.18)m in the corridor $(2.0m{\times}9.8m)$. To evaluate the effectiveness of the proposed method under rapid translational speed changes and sharp turns, we present the corresponding range of translational and rotational velocities, which indicate instances of aggressive motion. A total of six sequences are detailed in Table VII.
TABLE VII: Description of the sequences of the office dataset.
Seq
Duration
Length
Translational
velocity
max/min (m/s)
Rotational
velocity
max/min (rad/s)
Description
C-S1 526.4 51.01 0.15/0.10 0.20/0.12 Slow sequences
C-S2 474.2 53.24 0.17/0.12 0.20/0.15 Slow sequences
C-J1 138.6 37.42 0.26/0.22 1.74/0.20 Fast sequences
C-J2 159.8 37.30 0.26/0.22 1.76/0.22 Fast sequences
C-H1 242.5 39.95 0.26/0.13 1.74/0.12 Hybrid sequences
C-H2 198.6 36.36 0.26/0.12 1.75/0.14 Hybrid sequences
  • Slow sequences: the robot moves smoothly with low translational and rotational velocities, following straight and rectangular trajectories.
  • Fast sequences: the robot exhibits high translational and rotational velocities with rapid speed variations, following both a rectangular path and a wave path.
  • Hybrid sequences: the robot follows a combination of wave and rectangular trajectories, featuring alternating phases of slow and fast translational and rotational motion.
Fig. 11 shows the overview of trajectories in the office dataset. Fig. 12 shows the distribution of control points in sequence C-H1. The APE results of the compared methods on the office dataset are summarized in Table VIII.
According to the APE results, compared with the existing state-of-the-art discrete-time and continuous-time methods, our CT-UIO outperforms the existing methods and achieves the best results in most sequences. Furthermore, the estimated trajectories of our method are in good agreement with the ground truth in all the sequences. As for SFUISE, its localization results on C-S1 and C-S2 are close to CT-UIO. However, large errors are generated due to the fast translational and rotational motion in fast and hybrid sequences. Conversely, the proposed CT-UIO achieves consistent localization estimation. This can be attributed to the proposed adaptive knot span adjustment strategy that adjusts the density of control points based on the robot speed ranges during rapid and hybrid motion.
CDF plots are also an additional visual representation of the positioning error in Fig. 13. All the CDFs of our method are higher than those of the five state-of-the-art methods. For sequence C-H1, $90\%$ of the estimates from our method achieve APE less than $0.203m$, better than Optimization+Filtering ($0.974m$), Optimization+Trust Region ($0.324m$), GPDA ($0.261m$), VMHE ($0.269m$), and SFUISE ($0.231m$).
TABLE VIII: Performance comparison in APE (RMSE, Meter) of different methods on the office dataset. The best results are marked in bold.
Seq
C-S1
C-S2 C-J1 C-J2 C-H1 C-H2
Optimization+
Filtering[yang2023novel]
0.655 0.764 0.804 0.760 0.661 0.691
Optimization+
Trust Region[zhou2024optimization]
0.587 0.652 0.737 0.638 0.537 0.629
GPDA[WOS:001173317800046] 0.499 0.515 0.412 0.542 0.384 0.591
VMHE[9829196] 0.447 0.486 0.329 0.535 0.272 0.349
SFUISE[li2023continuous] 0.357 0.333 0.300 0.250 0.223 0.261
Ours 0.352 0.320 0.263 0.202 0.189 0.208
@@ C-S2 C-J1 C-J2 C-H1 C-H2
The results of these experiments in the corridor, exhibition hall, and office datasets prove that the proposed CT-UIO is highly effective in situations involving fast translational and rotational motion. When the UWB/IMU/odometer sensors are not time-synchronized, the discrete-time formulation (Optimization+Filtering, Optimization+Trust Region, GPDA, and VMHE) suffers from a loss in accuracy, particularly under fast motion, due to its reliance on a constant velocity assumption in the period of time between consecutive UWB ranging measurements. The hypothesis of time synchronization may not always hold for multiple sensors and fast motions. In contrast, the continuous-time formulation (SFUISE and our proposed CT-UIO) effectively handles asynchronous measurements and enables accurate trajectory estimates to be sampled at any time, thereby maintaining high positioning accuracy. Across all sequences, the CT-UIO maintains more stable and consistent accuracy by utilizing a non-uniform continuous-time trajectory representation. This strategy effectively adjusts the control point interval and ensures better capture of motion pattern characteristics.
V-E Ablation Study
To further evaluate the effectiveness of each novelty in the proposed method, ablation experiments are conducted on the corridor dataset, exhibition hall dataset, and office dataset. The APE results of robot localization under different ablation conditions are presented in Table IX. Ours (None) refers to our proposed method with the adaptive knot span adjustment strategy, improved EKF module, and VA generation disabled. The uniform B-splines and conventional EKF are utilized. Ours (AKSA), Ours (IEKF), and Ours (VA) denote variants of Ours (None) with the adaptive knot span adjustment, improved EKF module, and VA generation enabled individually, respectively. Ours represents the full implementation of our proposed method.
The effectiveness of the adaptive knot span adjustment strategy in handling dynamic motion is highlighted in the fast and hybrid sequences. By adjusting knot spans in response to sudden motion changes, the strategy enhances the localization robustness. As shown in Table IX, both Ours (IEKF) and Ours (VA) consistently outperform Ours (None) across all sequences. While IMU are prone to noise, particularly during aggressive motions, odometry provides more reliable velocity data, enabling effective sensor fusion. Ours (IEKF) constructs an IMU/odometer fusion model using an improved Extended Kalman Filter (EKF) with innovation-based adaptive estimation, successfully integrating IMU and odometry data. Ours (VA) further leverages the IMU/odometer fusion model to generate VAs, providing additional observation constraints for the UWB location system and enhancing overall system stability. When all modules are enabled, the full framework Ours achieves the highest overall accuracy and robustness.
TABLE IX: Performance comparison in APE (RMSE, Meter) under different conditions in ablation.
Method Corridor Dataset Exhibition Hall Dataset Office Dataset
A-S1 A-S2 A-J1 A-J2 A-H1 A-H2 B-S1 B-S2 B-J1 B-J2 B-H1 B-H2 C-S1 C-S2 C-J1 C-J2 C-H1 C-H2
Ours (None) 0.270 0.266 0.482 0.503 0.512 0.488 0.052 0.138 0.344 0.320 0.405 0.396 0.381 0.357 0.341 0.342 0.267 0.305
Ours (AKSA) 0.243 0.255 0.190 0.182 0.455 0.428 0.045 0.135 0.220 0.138 0.167 0.330 0.360 0.345 0.305 0.241 0.207 0.249
Ours (IEKF) 0.250 0.235 0.406 0.431 0.492 0.457 0.043 0.132 0.268 0.161 0.264 0.351 0.344 0.352 0.331 0.249 0.243 0.252
Ours (VA) 0.262 0.240 0.451 0.482 0.456 0.467 0.046 0.134 0.306 0.291 0.347 0.329 0.362 0.348 0.306 0.247 0.230 0.235
Ours 0.208 0.242 0.181 0.153 0.403 0.416 0.040 0.122 0.219 0.098 0.150 0.311 0.352 0.320 0.263 0.202 0.189 0.208
V-F Computational Costs
In this part, we investigate the time efficiency of continuous-time methods on the corridor dataset, and Table X shows the results. The time consumptions for optimization of CT-UIO are $12.64\mathrm{ms}$, $26.07\mathrm{ms}$, and $21.42\mathrm{ms}$ for the slow, fast, and hybrid sequences, respectively. In slow sequences, the motion is typically gentle, resulting in a larger knot span and fewer control points to be optimized, thereby reducing the optimization time. In contrast, in fast or hybrid sequences, larger variations in translational or rotational velocity require an increased number of control points to improve accuracy, which in turn leads to higher computational costs. For comparison, SFUISE employs a uniform control point insertion strategy, configured with $10$ control points at $1Hz$ as suggested in [li2023continuous], regardless of motion characteristics, leading to a more consistent but potentially less efficient time performance.
TABLE X: The average time consumption for optimization (msms).
Sequences Slow sequences Fast sequences Hybrid sequences
SFUISE[li2023continuous] 24.50 27.61 25.83
Ours 12.64 26.07 21.42
VI Conclusion
@@
In this article, we present CT-UIO, a continuous-time UWB-Inertial-odometer localization system designed to achieve accurate and robust localization under few-anchor conditions. CT-UIO tightly integrates UWB ranging, IMU, and odometer data using a non-uniform B-spline trajectory representation. Specifically, the proposed system enhances state estimation accuracy in fast-moving scenarios by adaptively placing control points according to varying motion speeds. To ensure the UWB-based localization system is fully observable, CT-UIO combines UWB ranging with motion prior from the IMU/odometer fusion model to generate VAs. Additionally, CT-UIO’s backend conducts adaptive sliding window factor pose graph optimization, incorporating constraints from VA’s ranging, IMU, and odometer measurements. The proposed CT-UIO system is evaluated in several real-world few-anchor situations using self-collected datasets. Compared to state-of-the-art discrete-time and continuous-time estimation methods, CT-UIO consistently produces more consistent and robust trajectory estimation, particularly in fast-motion scenes.
@@
In this work, the anchor positions are manually pre-calibrated using laser meters, which limits the applicability of the proposed CT-UIO method in UWB networks with unknown anchor locations or scenarios that require fully automated deployment without manual intervention. In the future, it is worthwhile exploring the implementation of the proposed CT-UIO method with unknown anchor positions, as well as incorporating complementary sensors, such as cameras, to enhance the accuracy of short-term pose estimates.